OpenStudy (anonymous):

Find all z∈ℂ (complex numbers) which satisfy (conjugate of z)=z^2

OpenStudy (jamesj):

Write z = a +bi. Then put that in the equation and find two equations in a and b. Solve.

OpenStudy (jamesj):

i.e., z conjugate, z* = a - bi; and z^2 = a^2 + 2abi + b^2i^2 = ... something. Now put z* = z^2. The real and imaginary parts must be equal, which will give you two equations in a and b.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

these repeats are coming fast and furious. where are you all from?

OpenStudy (anonymous):

i just did this one a moment ago

OpenStudy (anonymous):

yiazmat, bluebrandon also posting the same questions

OpenStudy (anonymous):

y'all should form a study group

OpenStudy (anonymous):

I already asked this question a little while ago lol

OpenStudy (jamesj):

@bluebrandon, where are you all from?

OpenStudy (anonymous):

so did yiazmat. where are you all from?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

university of waterloo

OpenStudy (jamesj):

and what class is this?

OpenStudy (anonymous):

this is for algebra

OpenStudy (turingtest):

fancy algebra in that university

OpenStudy (anonymous):

i'll say. complex variable algebra.

OpenStudy (jamesj):

found the answer?

OpenStudy (anonymous):

I got z=(-1+0i), (1+2i), (1-2i)

OpenStudy (jamesj):

Well, if z = -1, then z^2 = 1 and z* = -1, so that can't be right.

OpenStudy (jamesj):

You should verify your other solutions as well.

OpenStudy (anonymous):

woops wrong question for this one I got (x^2-y^2) (2xy) --------- - --------- i (x^2+y^2) (x^2+y^2)

OpenStudy (jamesj):

Also not right, I don't think.

OpenStudy (anonymous):

I put z* = z^2 (z)(z) = z* z= z*/z then I let z=(x+yi) and solved

OpenStudy (jamesj):

If z = a + bi, then z^2 = (a^2 - b^2) + 2abi and z* = a - bi, where * means conjugate. Hence if z^2 = z*, then (a^2 - b^2) + 2abi = a - bi ---- (**) Now for this to be true, the real and imaginary components must be equal. Hence (**) is true if and only if a^2 - b^2 = a ------ (1) 2ab = -b ------ (2)

OpenStudy (jamesj):

From equation (2), if b is not zero, then 2a = -1 => a = -1/2. Now solve for b. Notice that if b = 0, then you can show a = 0 or 1; i.e., z = 0 and z= 1 are solutions, which makes sense.

OpenStudy (anonymous):

does the way I've done it look like it could make sense?

OpenStudy (jamesj):

No, because you've written down an infinite family of solutions. But there are only a finite number.

OpenStudy (anonymous):

ok I will try to do it your way then

OpenStudy (anonymous):

how does z=0,1,(-1/2 + (sqrt3)/2), (-1/2 - (sqrt3)/2) sound?

OpenStudy (jamesj):

Yep. Those are the four solutions; just put i after the imaginary parts.

OpenStudy (anonymous):

alright, thanks for your help!